# Pedagogical Alignment of Large Language Models

### v1

@@ Abstract
In this paper, we introduce the novel concept of pedagogically aligned Large Language Models (LLMs) that signifies a transformative shift in the application of LLMs within educational contexts. Rather than providing direct responses to user queries, pedagogically-aligned LLMs function as scaffolding tools, breaking complex problems into manageable subproblems and guiding students towards the final answer through constructive feedback and hints. The objective is to equip learners with problem-solving strategies that deepen their understanding and internalization of the subject matter. Previous research in this field has primarily applied the supervised finetuning approach without framing the objective as an alignment problem, hence not employing reinforcement learning through human feedback (RLHF) methods. This study reinterprets the narrative by viewing the task through the lens of alignment and demonstrates how RLHF methods emerge naturally as a superior alternative for aligning LLM behaviour. Building on this perspective, we propose a novel approach for constructing a reward dataset specifically designed for the pedagogical alignment of LLMs. We apply three state-of-the-art RLHF algorithms and find that they outperform SFT significantly. Our qualitative analyses across model differences and hyperparameter sensitivity further validate the superiority of RLHF over SFT. Also, our study sheds light on the potential of online feedback for enhancing the performance of pedagogically-aligned LLMs, thus providing valuable insights for the advancement of these models in educational settings.
@@ 1 Introduction
Large Language Models (LLMs) are an integral part of the modern digital landscape, and ensuring their alignment to be honest, harmless, and helpful - the ‘3H’ principle - is critical to their effective usage [1]. However, the needs of a student learning a subject or preparing for an exam differ substantially from those of a casual internet user. While the latter might seek immediate answers to their queries, the former benefits more from a structured, guided approach that encourages active learning [2] and critical thinking [3]. Therefore, if we aim to leverage LLMs to assist students’ learning, we must redefine the ‘helpfulness’ criterion in a way that specifically supports this demographic [4].
@@
In this paper, we propose a novel concept of pedagogical alignment for LLMs which aligns LLMs to function as scaffolding tools [5], breaking complex problems into manageable subproblems and guiding students towards the final answer through constructive feedback and hint [6]. This approach redefines the ‘helpfulness’ criterion of the ‘3H’ principle, tailoring it to the unique needs and learning objectives of students. We argue that to truly assist students, LLMs should not merely provide immediate answers, but rather guide learners through a structured, step-by-step process that encourages active learning and critical thinking. The pedagogical alignment of LLMs is not just a theoretical construct; it is grounded in the belief that learning can be significantly improved when students are given the opportunity to actively participate in their learning process [7] guided by feedback [8].
Our work builds upon the framework of Conversational Learning with Analytical Step-by-Step Strategies (CLASS) [9, 10], a tool that applies above principles of learning sciences to construct an Intelligent Tutoring System (ITS) using LLMs. However, their approach is implemented through supervised fine-tuning of the LLM, which we identify as a limitation. We argue that the desired behavior to foster learning can be more effectively modeled using reinforcement learning. However, the challenge lies in the absence of a reward dataset necessary for reinforcement learning in education context.
To address this challenge, we reinterpret the CLASS framework from a pedagogical alignment perspective, proposing that it can be viewed as a potential reinforcement learning problem where the rewards are associated with pedagogical alignment. This perspective allows us to construct a preference (reward) dataset, laying the groundwork to enhance the alignment of LLMs to better serve educational purposes.
@@
Next, using this reward dataset, we study the effective of three RLHF algorithms: Direct Preference Optimization (DPO) [11], Identity Preference Optimization (IPO) [12], and Kahneman Tversky Optimization (KTO) [13] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [14] and Zephyr [15]. We measure the the efficacy of these algorithms on pedagogical alignment by an accuracy metric. This metric quantifies the alignment by considering whether the correct pedagogical action was taken by the model - a crucial aspect of our pedagogical alignment approach.
Our findings reveal that DPO and KTO significantly enhance the performance of both Mistral and Zephyr models, improving our pedagogical alignment accuracy by 37 and 23 points respectively, compared to CLASS’ SFT approach. Also, we examine the influence of the beta hyperparameter on the performance of these alignment algorithms since beta is a critical parameter that significantly affects the outcome of the alignment process.
Our contributions in this paper are threefold:
  1. 1.
    We introduce the concept of “pedagogical alignment,” a new approach to align LLMs in a manner that fosters optimal student learning.
  2. 2.
    We reinterpret the CLASS framework from a pedagogical alignment perspective, which allows us to create a reward dataset.
  3. 3.
    Using the reward dataset, we demonstrate the effectiveness of reinforcement learning-based alignment algorithms on state-of-the-art LLMs.
Through this work, we aim to contribute to the effective use of LLMs in education, highlighting their potential as productive learning aids and enhancing the learning experience for students.
Table 1: Inference prompt of CLASS framework.
@@
Prompt: Instructions to act as a Tutorbot, an AI-powered chatbot designed to help student with a question. Your goal is to break problem into  subproblems for the student, and walk the student to the final answer using the subproblems and hints.
For each student response, think which category your response falls in, and then frame the reply using this json structure:
{ "Evaluation": "a,b,c,d,e,f,g",
   "Action Based on Evaluation": "1,2,3,4,5,6,..,12",
   "Subproblem State": "x,y,z",
   "Subproblem": "..",
   "Tutorbot": "..", }
@@
a) Correct Response
b) Incorrect Response
c) Partially Correct Response
d) Ambiguous or Unclear Response
e) ...
Tutorbot Actions:
@@
1) Confirm the correct answer. Check for completeness of the answer. If solution is incomplete, notify the student to complete the solution.
Prompt continues: If "b" is the evaluation, then: 2) Notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer hints 3) Step in to provide a solution if the student is unable to answer even after multiple attempts. If "c" is the evaluation, then: 4) Acknowledge the accurate parts, Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints 5) Step in to provide a solution if the student is unable to answer after multiple attempts If "d" is the evaluation, then: 6) Actively seek clarification through relevant follow-up questions. Request student to provide specific information Subproblem State: x) One of the subproblems is currently being solved y) Subproblem finished, move to next subproblem z) Subproblem finished, no next subproblem, problem finished
2 CLASS through the lens of pedagogical alignment
This section presents the concept of pedagogical alignment and uses it to examine the CLASS framework. This examination highlights the potential of application of reinforcement learning and paves the way for constructing a reward dataset to enhance the pedagogical alignment of LLMs.
2.1 Pedagogical alignment
In the landscape of LLMs, the term “alignment" is a concept used to describe the process of adapting LLMs to align with human preferences. In the context of online conversational agents, alignment specifically refers to making an LLM honest, harmless, and helpful to a typical internet user [1]. However, the needs of a student are distinct from those of a regular internet user. As illustrated in Figure 1, when an internet user poses a question to a LLM, the LLM’s response should directly address the user’s query. This is the primary use case for out-of-the-box LLMs like Mistral [14] and Llama [16], which are designed for such interactions. However, to promote optimal student learning, LLMs should guide the student to reach the correct answer through their own reasoning by employing effective scaffolding [17], involving problem decomposition, constructive feedback, and timely hints.
Here, we introduce our unique perspective of viewing the CLASS framework in the context of pedagogical alignment. The CLASS tutorbot differentiates itself from conventional out-of-the-box LLMs through its approach to student queries. Instead of directly providing answers, the CLASS tutorbot takes on a more interactive and guided approach. It begins by breaking the problem into smaller, manageable subproblems. Each subproblem presents a learning opportunity for the student, guiding them towards the final solution in a step-by-step manner.
The CLASS framework achieves effective scaffolding through a defined discrete action and state space. For every student response, the CLASS tutorbot determines whether the student’s response is correct, incorrect, or ambiguous. For instance, if the response is incorrect, an out-of-the-box LLM’s instinctual response would be to provide the correct answer to the user. However, the CLASS tutorbot’s initial reaction is to offer constructive feedback to the student and allow the student to arrive at the right answer independently. If the student is unable to do so, then the CLASS’s eventual response is to step in and provide the correct answer (refer to action 2 and 3 in inference prompt  1).
Similarly, if the student’s response is partially correct, the CLASS tutorbot’s first reaction is to inform the student about the correct aspects and then highlight the errors, encouraging the student to correct their answer independently (action 4 in prompt 1), as opposed to providing the complete answer, which is a typical reaction of an out-of-the-box LLM.
Pedagogical alignment: We define the term ‘pedagogical alignment’ to describe this shift in behavior that CLASS introduces, in contrast to an out-of-the-box LLM, to promote learning [18]. While out-of-the-box LLMs are designed to be helpful to general users, pedagogically aligned LLMs are tailored to foster optimal learning experiences for students.
2.2 Reinforcement learning algorithms for alignment
@@
While the CLASS framework used SFT for pedagogical alignment, recent developments in the field of Reinforcement Learning from Human Feedback (RLHF) present promising alternatives. RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives. The seminal work by [19] introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF. Subsequent advancements [20, 21] have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
The RLHF methodology unfolds in a structured three-step process:
  1. 1.
    Supervised Fine-Tuning (SFT): Initially, the LLM undergoes SFT using a task-specific dataset. This step adjusts the model’s initial parameters to align with the downstream task’s objectives, providing a tailored foundation for further refinement.
  2. 2.
    Preference Modeling: Following SFT, the model generates multiple responses under varied generation parameters. Human evaluators then rank these responses based on their preferences. This ranking process informs the development of a preference model, which learns to assign scores to LLM outputs, effectively translating human judgments into quantifiable metrics.
  3. 3.
    Online Reinforcement Learning: Utilizing the preference model, LLMs are further aligned through online reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [22]. This phase iteratively refines the model’s parameters to maximize the alignment with human feedback, enhancing the model’s performance and ethical alignment.
@@
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (1)
where $\sigma$ is the sigmoid function, and $\beta$ is an algorithmic parameter. However, DPO has been observed to over-fit to the preference dataset, and is believed to be sensitive to hyperparameter adjustments. To prevent the over-fitting behaviour of DPO’s optimization algorithm, recent proposals include Identity Preference Optimisation (IPO) [12] and Kahneman-Tversky Optimization (KTO) [13].
All the three methodologies, DPO, IPO, and KTO, require a dataset composed of preferences, denoted as ${(x,y_{w},y_{l})}$. The assembly of such datasets in a real-world context is notably labor-intensive and financially burdensome. However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for efficient and cost-effective pedagogical alignment of LLMs.
@@ 3 Preference Data Generation for Pedagogical Alignment
Figure 2: This illustration shows the process of generating pedagogically-aligned preference data using the CLASS framework. The GPT-student asks a question, to which both the GPT-tutor and SFT-tutor respond. The key focus here is on the divergence in the ‘Action Based on Evaluation’ between the two tutors. In this example, the GPT-tutor’s response is deemed more pedagogically aligned because it encourages the student to engage in critical thinking and attempt the problem again, instead of directly providing the correct answer. This action mismatch between the tutor responses allows us to construct a preference dataset that distinguishes between the pedagogically preferred (chosen) and less effective (rejected) responses. The preference dataset serves as the foundation for applying reinforcement learning-based alignment algorithms, with the goal of enhancing the pedagogical alignment of large language models.
In this section, we describe our DPO, IPO, and KTO compatible data generation pipeline. The objective is to generate a preference dataset $\mathcal{D}_{p}$ which necessitates distinguishing between pedagogically chosen and rejected tutor responses based on their alignment with desired scaffolding strategies. $\mathcal{D}_{p}$ will mimic a dataset structure that is compatible with DPO, IPO, and KTO to optimally represent these preferences. This dataset has three fundamental components: context, chosen, and rejected. In our case, the context represents the accumulated conversational history between the student and the tutor leading up to a particular interaction point, providing the necessary backdrop for the tutor’s next dialogue turn. The chosen contains the pedagogically preferred or optimal tutor responses to the context, while rejected includes those tutor responses considered less effective or misaligned with the scaffolding approach.
CLASS Conversational Dataset $\mathcal{D}_{c}$ and Agents: Our base dataset is the CLASS conversational dataset represented by $\mathcal{D}_{c}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,\ldots,N\}$, comprising $N$ student-tutor conversational turns. Each $\mathbf{x}$ represents a sequence of student utterances, and each corresponding $\mathbf{y}$ represents the tutor response.
@@
$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Evaluation of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Evaluation": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
@@
$\text{and }f(\mathbf{y}_{t}^{\mathcal{A}_{T}})\neq f(\mathbf{y}_{t}^{\mathcal{A}_{S}}),\forall t\in\{1,\ldots,N\}\big{\}}$ (4)
where $C_{t}=x,y_{w}=y_{choosen},y_{l}=y_{rejected}$ from DPO equation (1) from section 2.2, and the function $f$ is the key to distinguishing between chosen and rejected tutor responses. $f$ indicates a discrepancy between $\mathcal{A}_{\text{T}}$’s and $\mathcal{A}_{\text{S}}$’s responses if any of the fields “Evaluation", “Action Based on Evaluation", or “Subproblem State" diverge between $\mathcal{A}_{\text{T}}$ and $\mathcal{A}_{\text{S}}$, $f$ asserts $\mathcal{A}_{\text{T}}$’s response as preferable, under the assumption that $\mathcal{A}_{\text{T}}$ offers a more accurate pedagogical model, alongside the rejected response from $\mathcal{A}_{\text{S}}$. This condition is captured through the function $f$—mapping responses to their pedagogical attributes:
$f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\,\mathrm{ActionEval}(\mathbf{y}_{i}^{\mathcal{A}}),\,\mathrm{SubprobState}(\mathbf{y}_{i}^{\mathcal{A}})\right)$ (5)
where,
@@
  • $\text{ActionEval}(\mathbf{y}_{i}^{\mathcal{A}})$ outlines the tutor’s recommended actions based on the evaluation.
  • $\text{SubprobState}(\mathbf{y}_{i}^{\mathcal{A}})$ provides an understanding of the tutor’s perception of the student’s current understanding or progress on sub-problems within a larger problem context.

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable potential in various domains, including education. However, their application in educational contexts requires careful ‘pedagogical alignment’, where these models are tuned to assist students in a step-by-step problem-solving process, rather than providing direct answers. In this paper, we investigate algorithms for learning from human preferences (LHP) to achieve this alignment objective. A key challenge in this process is the lack of preference datasets to guide the alignment. To address this, we propose a novel approach for constructing such a dataset using synthetic data generation techniques. Leveraging this dataset, our experiments reveal the superiority of LHP methods over standard supervised fine-tuning (SFT) for pedagogical alignment. Applying LHP algorithms to Mistral and Zephyr, we observe a significant improvement in their ability to guide students, enhancing the pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT. Additionally, we introduce a unique approach for evaluating the pedagogical effectiveness of LLMs using perplexity analysis. We curate a second preference dataset, comparing the perplexity of LLMs’ responses when offering scaffolded step-by-step guidance versus direct solutions. This allows us to quantify the extent to which LLMs prioritize pedagogically effective actions in their generation. Our analysis provides compelling evidence for the superiority of LHP methods over SFT in optimizing LLMs’ behavior, underscoring the potential of LHP methods in better aligning LLMs with educational objectives and fostering effective learning experiences.
@@ 1 Introduction
As Large Language Models (LLMs) become integral to the digital landscape, aligning them with the ‘3H’ principle - honesty, harmlessness, and helpfulness - is crucial [1]. In an educational context, ‘helpfulness’ takes on a new meaning; rather than providing immediate answers, LLMs should guide students through a structured step-by-step problem-solving processes to encourage active learning and critical thinking [2, 3, 4]. This objective, which we term ‘pedagogical alignment,’ involves decomposing complex problems and providing constructive feedback and hints [5, 6]. Previous adaptation of LLMs for education, such as the CLASS framework [7], relied on supervised fine-tuning (SFT). However, SFT doesn’t explicitly favor pedagogically effective responses, a key aspect of pedagogical alignment [8, 9].
@@
In this paper, we address this limitation by modeling pedagogical alignment as a learning from human preferences (LHP) problem, where desired alignment is represented by preferences. A major challenge in this approach is the absence of a preference dataset tailored for the educational context. To overcome this, we devise a method for generating a synthetic preference dataset, extending the conversational dataset generation pipeline used in CLASS [7].
@@
Next, using this preference dataset, we study the effectiveness of three LHP algorithms: Direct Preference Optimization (DPO) [10], Identity Preference Optimization (IPO) [11], and Kahneman Tversky Optimization (KTO) [12] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [13] and Zephyr [14]. We evaluate the LLMs using pedagogical alignment accuracy and F1 score which measure if the models produce the desired pedagogical responses. Our results show that DPO and KTO significantly improve Mistral and Zephyr’s performance, improving pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT.
To further investigate the effectiveness of LHP over SFT, we introduce a second preference dataset generation technique. This technique allows us to conduct a perplexity analysis, examining the LLM’s predisposition to provide hints and guidance versus giving direct answers. This analysis is critical in understanding how well the LLM aligns with our definition of pedagogical alignment, which emphasizes providing step-by-step guidance over direct answers. We used this technique to generate pedagogically misaligned actions, which we then compared to the aligned actions in terms of perplexity. Our analysis revealed that base models are more inclined to provide direct solutions than offer hints and guidance. While SFT partially corrects this behavior, it doesn’t fully optimize it. However, LHP dramatically optimizes this behavior, promoting guidance over direct solutions. This shift in behavior underscores the effectiveness of LHP in aligning LLMs with educational objectives, providing compelling evidence for its superiority over SFT in fostering effective learning experiences.
2 Related Work
@@
Prompt: Your goal is to create a mock conversation between Student and a Tutorbot, an AI-powered chatbot designed to help Students with a question:
"Student": "Q. {problem}",
"Evaluation of Student Response": "..",
"Action Based on Evaluation": "..",
"Subproblem State": "..",
"Tutorbot": "Lets break the problem into subproblems and tackle the subproblems one by one. Lets begin with the first subproblem...",
@@
a) Evaluating Incorrect Responses
b) Evaluating Correct Responses
c) Evaluating Partially Correct Responses
d) Evaluating Ambiguous or Unclear or Short Responses
e) Redirecting Off-topic Responses
f) Responding to Student Inquiries
g) N/A
Prompt continues:
@@
Action 1) Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints
Action 2) Step in to provide a solution if the student is unable to answer even after multiple attempts.
If "c" is the evaluation, then:
Action 4) Acknowledge the accurate parts, Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints
Action 5) Step in to provide a solution if the student is unable to answer after multiple attempts
Subproblem states=:
x) One of the subproblems is currently being solved
y) Subproblem finished, moving to next subproblem
Table 1: Synthetic conversational data generation prompt of CLASS framework. While CLASS introduced actions for interpretability, in our work we utilize these actions in a novel way to construct a preference dataset that distinguishes pedagogically preferred responses from less effective ones for training our LHP models.
2.1 Algorithms for LHP
@@
Recent developments in the field of learning from human preferences (LHP) and reinforcement learning through human feedback (RLHF) present promising alternatives for pedagogical alignment. LHP/RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives. The seminal work by [9] introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF. Subsequent advancements [15, 16] have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
@@
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
where $\sigma$ is the sigmoid function, and $\beta$ is an algorithmic parameter.
However, DPO has been observed to over-fit to the preference dataset, and is believed to be sensitive to hyperparameter adjustments. To prevent the over-fitting behaviour of DPO’s optimization algorithm, recent proposals include Identity Preference Optimisation (IPO) [11] and Kahneman-Tversky Optimization (KTO) [12].
All the three methodologies, DPO, IPO, and KTO, require a dataset composed of preferences, denoted as ${(x,y_{w},y_{l})}$. The curation of such datasets in a real-world context is notably labor-intensive and financially burdensome. However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for cost-effective pedagogical alignment of LLMs.
2.2 Synthetic Student Data Generation
To create our preference dataset, we adapt the CLASS framework [7] that uses GPT-4 [17] to generate synthetic conversational datasets. These datasets simulate a range of interactions between students and GPT-4 using various prompts. An example prompt is illustrated in Table 1. The prompt instructs the model to first evaluate the student response (Evaluation of Student Response), consider which action to employ (Action based on Evaluation), determine the sub-problem (Subproblem State), and conditioned on these fields generate the reply to the student. While the CLASS framework introduced these fields to enhance interpretability for teachers and other educational stakeholders, we utilize them differently to construct a preference dataset, as we will discuss in the next section.
@@ 3 Preference Data Generation for Pedagogical Alignment
Figure 2: This figure shows the process of generating pedagogically-aligned preference data using the CLASS framework. The GPT-student asks a question, to which both the GPT-tutor and SFT-tutor respond. The key focus here is on the divergence in the ‘Action Based on Evaluation’ between the two tutors. In this example, the GPT-tutor’s response is deemed more pedagogically aligned because it encourages the student to engage in critical thinking and attempt the problem again, instead of directly providing the correct answer. This action mismatch between the two tutor responses allows us to construct a preference dataset that distinguishes between the pedagogically preferred (chosen) and less effective (rejected) responses.
In this section, we describe our DPO, IPO, and KTO compatible data generation pipeline which we also outline in figure 2. The objective is to generate a preference dataset $\mathcal{D}_{p}$ which necessitates distinguishing between pedagogically chosen and rejected tutor responses based on their alignment with desired scaffolding strategies. $\mathcal{D}_{p}$ will mimic a dataset structure that is compatible with DPO, IPO, and KTO to optimally represent these preferences. This dataset has three fundamental components: context, chosen, and rejected. In our case, the context represents the accumulated conversational history between the student and the tutor leading up to a particular interaction point, providing the necessary backdrop for the tutor’s next dialogue turn. The chosen contains the pedagogically preferred or optimal tutor responses to the context, while rejected includes those tutor responses considered less effective or misaligned with the scaffolding approach.
Conversational Dataset $\mathcal{D}_{c}$ and Agents: We use the conversational prompt listed in Table 1 to generate a dataset represented by $\mathcal{D}_{c}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,\ldots,N\}$, comprising $N$ student-tutor conversational turns. Each $\mathbf{x}$ represents a sequence of student utterances, and each corresponding $\mathbf{y}$ represents the tutor response.
@@
$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Eval of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Eval": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
@@
where $C_{t}=x,y_{w}=y_{choosen},y_{l}=y_{rejected}$ from DPO equation (1) from section 2.1, and the function $f$ is the key to distinguishing between chosen and rejected tutor responses. $f$ indicates a discrepancy between $\mathcal{A}_{\text{T}}$’s and $\mathcal{A}_{\text{S}}$’s responses if any of the fields “Evaluation", “Action Based on Evaluation", or “Subproblem State" diverge between $\mathcal{A}_{\text{T}}$ and $\mathcal{A}_{\text{S}}$, $f$ asserts $\mathcal{A}_{\text{T}}$’s response as preferable, under the assumption that $\mathcal{A}_{\text{T}}$ offers a more accurate pedagogical model, alongside the rejected response from $\mathcal{A}_{\text{S}}$. Figure 2 illustrates this conception with a concrete example: GPT-tutor ($\mathcal{A}_{\text{T}}$) and SFT-tutor ($\mathcal{A}_{\text{S}}$) provide different responses since the Action fields are different, so we choose the GPT-tutor response as the accepted one and the SFT-tutor response as the rejected one. This condition is captured through the function $f$—mapping responses to their pedagogical attributes:
${f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Action}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})\right)}$ (5)
@@
  • $\text{Action}(\mathbf{y}_{i}^{\mathcal{A}})$ outlines the tutor’s recommended actions based on the evaluation.
  • $\text{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})$ provides an understanding of the tutor’s perception of the student’s current understanding or progress on sub-problems within a larger problem context.